OpenStudy (anonymous):

Question: "Which of the following is equivalent to the equation 20-4/5x is equal to OR less than 16?" >I've put 20-4/5x = 19 & 1/5x >So 19 & 1/5x is equal to or less than 16 is what I've gotten so far... help me get farther or if I'm doing this wrong? > The Answer IS A. x is equal to or less than 5 but don't know how to get there. Thanks!!

OpenStudy (johnweldon1993):

So you start out with \[20 - \frac{ 4 }{ 5 }x \le 16\] you want to isolate 'x' so subtract 20 from both sides \[ - \frac{ 4 }{ 5 }x \le -4\] Now you want to multiply both sides of the equation by -5/4 (the reciprocal (remember we talked about that? this will isolate 'x' \[ - \frac{ 4 }{ 5 }x * -\frac{5}{4}\le -4 * -\frac{5}{4}\] \[ x \ge -4 * -\frac{5}{4}\] **(now you just multiplied by a negative number so you SWITCH the sign as shown) so now...what is -4 times = 5? 20 right? so what is 20 / 4? it is 5 so we have \[x \ge 5\] So I get the answer as being x is GREATER THAN or equal to 5

OpenStudy (johnweldon1993):

typo up there "what is -4 times = 5" should be what is -4 times - 5 other than that...all good :)

OpenStudy (anonymous):

LOL I caught that :P Thanks...ONCE again!! :)

OpenStudy (johnweldon1993):

haha not a problem :) and hey..1 mistake ...not bad :P

OpenStudy (anonymous):

Not bad at all :) @johnweldon1993

OpenStudy (anonymous):

@johnweldon1993 & your probably be right...cause I looked and YOUR answer is B. so I'll double check the answer sheet.

OpenStudy (johnweldon1993):

Okay :) I'm just PRETTYYYY sure that is the correct answer :)